xy-1= (factoring)

Answer by troyapplehelen(46)

1 is always x(1/x) or y(1/y)
so, xy=x(1/x). then, xy=x(1/x)
in factoring, xy-x(1/x)=0 x(y-1/x)=0 answers x=1/y or y=1/x